Please help me with a step by step process for solving linear equations. My problem is:
x + y = 4
x - y = 5
I don't know what the first step should be.

Answer by spacewolf(2)



The very first step is to choose a variable to solve for. Traditionally, I always solve for "y" first, and then I'll solve for "x". Here's the problem, step by step:


x + y = 4x - y = 5 {Here, I'll solve for "y"}


x + y = 5 {Remember that the problem establishes that each part of the problem containing variables are equal to "5"}


y = 5 - x


4x - y = 5 {Again, each part containing variables is equal to "5"}


4x - (5 - x) = 5 {We solved for "y" earlier, so now we plug it in to the equation to solve for "x"}


5x - 5 = 5


5x = 10


x = 2


y = 5 - x {Go back to how we solved for "y" and plug in the value for "x"}


y = 5 - 2


y = 3


Presto: x = 2, and y = 3. Problem solved! The important step is to isolate one of the variables first and solve for it. Then you can plug and chug your way to answer. Hope this helps.